OpenStudy (anonymous):

Solve the equation. x - 9 = -6x + 5 A. 14 B. 7 C. 2 D. 21

OpenStudy (anonymous):

take -6x to the left of the equation

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

and -9 to the right

OpenStudy (uri):

x - 9 = -6x + 5 Add 9 both sides x - 9 = -6x + 5 +9 +9 ========== x=-6x+14 x+6x=14 7x=14 divide 14 with 7 x=14/7 x=2 ans.